# The mantle-inner core gravitational mode of oscillation in a strong magnetic field regime

### v1

@@
Abstract The mantle-inner core gravitational (MICG) mode is the free mode axial oscillation between the mantle and inner core sustained by the gravitational torque between their degree 2 order 2 density structures. Here, we investigate how the MICG mode is affected by oscillations of cylindrical surfaces in the fluid outer core in the form of Alfvén waves. The latter are triggered by oscillations of the tangent cylinder (TC) moving jointly with the inner core and propagate away from the rotation axis. We show that the MICG mode remains a distinct normal mode of oscillation of the core-mantle system only when the triggered Alfvén waves are attenuated before they traverse the width of the fluid core. For an internal magnetic field strength of a few mT, as we expect in Earth’s core, Alfvén waves can readily traverse the width of the core, and the MICG mode is absorbed into the spectrum of torsional oscillation (TO) modes. The MICG period retains a dynamical influence, acting as a point of resonance for TO modes, and marking the transition from a TO mode in which the motion of the TC (including the inner core) is weakly impacted by gravitational coupling to one in which the oscillating motion of the TC is strongly restricted. Our results imply that the observed 6-year periodic signal in the length of day cannot be interpreted as the signature of the MICG mode and must instead be caused by TO modes, or more generally, by the propagation of Alfvén waves. {summary}Non-technical summary The Earth’s mantle is not uniform but contains continental size density anomalies. The topography of the solid inner core (located at the centre of our planet) is aligned on average with these mantle density anomalies. The axial gravitational restoring torque between these mantle and inner core structures – when they are rotated out of longitudinal alignment – sustains a natural mode of oscillation termed the mantle-inner core gravitational (MICG) mode. Previous studies have suggested that the MICG mode may explain a periodic signal of 6 years observed in the changes in the length of day (LOD). Here, we show that when the magnetic field inside the fluid core is high (of the order of a few mT), the nature of the MICG changes; it is no longer a distinct, independent mode of oscillation of the core-mantle system. Instead, the MICG mode merges into the set of modes of torsional oscillations (TO), which consist in the longitudinal oscillations of nested cylindrical surfaces within the fluid core maintained by the magnetic tension between them (oscillations that also known as Alfvén waves). The inner core can still oscillate with respect to the mantle, but it is due to its entrainment as part of the TO modes. The MICG mode retains a dynamical influence, as its period determines whether the inner core oscillation for a given TO mode is impeded or not by the mantle gravitational torque. These results have important implications for the interpretation of the 6 year LOD signal, which cannot be caused by the MICG mode, but must instead be caused by the TO modes, or more generally, by the propagation of Alfvén waves in the fluid core.
@@ 1 Introduction
The mantle-inner core gravitational (MICG) mode of axial oscillation owes its existence to aspherical mass anomalies within the mantle, including those caused by the topography at surfaces of density contrast (buffett96a). The largest of these occur at spherical harmonic degree 2 order 2 (e.g. defraigne96; simmons07), warping equipotential surfaces everywhere inside the Earth (including the surface geoid) into ellipsoid shapes elongated in the equatorial plane. Within the core, surfaces of constant density coincide with these elliptically shaped equipotential surfaces in order to maintain hydrostatic equilibrium. Provided the inner core deforms viscously on a timescale shorter than the typical changes in the gravitational potential imposed by the mantle, the inner core boundary (ICB) should also match an equipotential surface. Its shape should then be elliptical in the plane of the equator (i.e. equatorially elliptical) and longitudinally aligned with the leading order mantle mass anomalies (buffett96a). An axial rotation of the inner core out of this longitudinal alignment, driven for example by magnetic stresses at the ICB (e.g. buffett00; aubert11), leads to a gravitational torque acting to restore the equilibrium and state of minimum gravitational potential energy (buffett96a). This gravitational torque can thus sustain a natural mode of axial oscillation between the inner core and the mantle, referred to as the MICG mode (mound06).
@@
The period of the MICG mode depends on the gravitational strength factor, $\overline{\Gamma}$, capturing the amplitude of the gravitational torque for a given small longitudinal misalignment between the inner core and mantle (buffett96b). In turn, $\overline{\Gamma}$ depends on the amplitude of the degree 2 order 2 geoid at the core-mantle boundary (CMB), which itself depends on estimates of the mantle density field, including that from the topography of the CMB (buffett96a; davies14; chao17; shih21). The period of the MICG mode also depends on whether the region of the fluid core inside the axial cylinder that encloses the inner core – the so-called tangent cylinder (TC) – is entrained into co-rotation with the inner core (see eqs. (8) and (9) below).
@@
Estimates of the period of the MICG mode can be derived by computing predictions of $\overline{\Gamma}$ on the basis of viscous mantle flow models. The latter are built from estimates of the density anomalies within the mantle inferred from seismic tomography and the subsequent viscous deformation of the CMB, the surface and other surfaces of density discontinuities that they cause (e.g. hager85; forte94; defraigne96). Based on the mantle flow models available at the time, buffett96a estimated a MICG period of approximately 2 to 3 years. The more recent survey of davies14 suggests a longer period in the range 7 to 18 years. The range of uncertainty reflects our limited knowledge of both the amplitude of the internal mantle density anomalies (primarily those of degree 2, order 2 associated with Large Low Seismic Velocity provinces in the lower mantle, e.g. garnero16; mcnamara19) and the viscosity of the lowermost mantle.
@@
An alternate approach is to seek the signature of the MICG mode in the observed changes in the mantle rotation rate measured as changes in the length of day (LOD). Indeed, an observed quasi-periodic LOD signal with an amplitude of $\sim 0.13$ ms and a period of $\sim 5.9$ days (abarcadelrio00; holme13; chao14) has been linked with the MICG mode (mound06; chao17; ding18; duan18; shih21). The magnitude of $\overline{\Gamma}$ required to match this period provides then a direct measure of the equatorial ellipticities of the geoid at the CMB and of the ICB topography (davies14; chao17; shih21).
@@
This simplified picture of the MICG mode neglects the dynamical influence of the fluid core in the region outside the tangent cylinder (ROTC) or assumes that it plays a passive role. At periods relevant to the MICG, axisymmetric azimuthal (zonal) flows are expected to be invariant in the direction of the rotation axis and to take the form of rigidly rotating nested cylindrical surfaces (jault08). A differential motion between adjacent cylinders shears the cylindrically radial ($B_{s}$) magnetic field lines that threads their surfaces. This creates an azimuthal restoring force between cylinders and allows for the propagation of Alfvén waves propagating in the cylindrically radial direction (braginsky70).
The normal modes of axial oscillations between the cylinders formed by the superposition of these Alfvén waves are referred to as torsional oscillations (TO, braginsky70) and their periods depend the strength of $B_{s}$ within the core. The prevalent view at the time the earliest studies on the MICG mode were published (e.g. buffett96a; buffett96b; mound06) was that the gravest mode of TO had a period of 60-80 years, which then implied a typical $B_{s}$-field strength of 0.2-0.3 mT within the core (zatman97). Under this scenario, a subdecadal MICG mode has a distinctly shorter period, so the ROTC may be considered to play a passive role even if it is engaged in the MICG to some degree by the propagation of Alfvén waves (e.g. mound03; mound06).
@@
The general consensus on the magnetic field strength within the core has since shifted. Numerical models of the dynamo suggest that the field deep inside the core is likely stronger by a factor 10 than that at the CMB, suggesting an internal field strength of 3-4 mT or higher (e.g. christensen06; aubert09). At such a strength, Alfvén waves should propagate through the entire width of the core in only a few years and the periods of the gravest TO modes are expected to be in the subdecadal range. Indeed, gillet10 showed that the angular momentum carried by propagating Alfvén waves in the ROTC matches that required to explain the observed periodic $5.9$-yr LOD signal.
This latter result indicates that, for the 5.9 yr LOD signal, the mantle exchanges its angular momentum primarily with the ROTC. This is in contrast to the oscillation behaviour expected for the MICG mode, where the mantle exchanges its angular momentum primarily with the inner core (whether entraining the entire TC or not) and not the ROTC. It also presents an alternative explanation for the 5.9 yr LOD signal, which would then represents the entrainment of the mantle into oscillation as a result of these propagating Alfvén waves. Or, at the very least, if the MICG mode remains connected to the 5.9 yr LOD signal, the ROTC is not passive and is instead entrained into motion by the propagation of Alfvén waves.
@@
Here, our goal is to investigate how the MICG mode is altered when a strong $B_{s}$-field inside the core allows for the propagation of fast Alfvén waves that can travel through the width of the core in a timescale shorter or similar to the MICG period. We shall show that the MICG mode is no longer an independent oscillation mode of the coupled core-mantle system but rather gets absorbed into the spectrum of TO modes, although the latter may be modified by gravitational coupling.
@@ 2.1 Angular momentum balance
The model of internal coupling between the mantle, fluid core and inner core that we use follows broadly that presented in several studies, including in buffett96b; buffett98; mound03; mound05b; dumberry08c; dumberry10b. We assume a spherical model of Earth’s interior comprised of an inner core (radius $r_{i}$), a fluid core (outer radius $r_{f}$) and mantle. We adopt a simplified model of flow in the fluid core restricted to its azimuthal component $v_{\phi}$ and assumed to be axisymmetric and invariant along the direction of the rotation axis. Accordingly, we use cylindrical coordinates ($s$, $\phi$, $z$), and we divide the fluid core in a set of $N_{s}$ discrete cylindrical surfaces with radius $s$ that intersect the CMB at axial position $z_{f}=\sqrt{r_{f}^{2}-s^{2}}$ in the Northern hemisphere. Cylinders inside the TC ($s<r_{i}$) intersect the ICB at $z_{i}=\sqrt{r_{i}^{2}-s^{2}}$; we set $z_{i}=0$ for cylinders outside the TC. The azimuthal velocity of each cylinder is set equal to $v_{\phi}=s\Omega_{f}$, where $\Omega_{f}\equiv\Omega_{f}(s)$ is the angular velocity.
@@
is the axial moment of inertia density of cylinders within the fluid core, where $\rho$ is the density (assumed uniform). $\Gamma_{cmb}$ and $\Gamma_{icb}$ are the torques from all surface forces acting on the mantle at the CMB and on the inner core at the ICB, respectively. $\overline{\Gamma}$ is the strength of gravitational torque between the mantle and inner core and $\alpha$ is the (assumed small) longitudinal angle of misalignment between the degree 2 order 2 mantle density field and the long equatorial axis of the ICB topography. $\alpha$ depends on the differential rotation between the inner core and the mantle and on the viscous relaxation time ${\tau}$ for the ICB topography to return to its equilibrium alignment with the imposed gravitational potential from the mantle; its evolution is given by
@@
We assume that $\Gamma_{icb}$ and $\Gamma_{cmb}$ are both dominated by electromagnetic (EM) forces, and expressed as (e.g. buffett98)
[…]
@@
where $B_{m}^{d}$ and $B_{i}^{d}$ are the polar amplitudes of the axial dipole at the CMB and ICB, respectively, $\langle B_{r,m}\rangle$ is the root mean squared (rms) strengths of the total radial field (including the axial dipole) at the CMB, $G_{m}$ is the conductance of the lowermost mantle and $G_{i}$ the conductance factor at the ICB given by
@@
where $\sigma$ is the conductivity of the fluid core (assumed equal to that of the inner core), $\delta=\sqrt{2/|\omega|\mu\sigma}$ is the magnetic skin depth at oscillation frequency $\omega$, with $\mu$ the magnetic permeability of free space, and $\mbox{sgn}(\omega)=\omega/|\omega|$. The form of ${\cal F}_{m}(s)$ was shown in dumberry08c to be a good match with the EM coupling resulting from a realistic field model at the CMB comprised of a large number of spherical harmonic terms. The field at the ICB is undoubtedly more complex than a simple axial dipole, although the latter is expected to be its dominant contribution, and it is sufficient to capture the first order effect of EM coupling at the ICB.
@@ 2.2 The MICG mode
To illustrate how the MICG mode emerges from the above system, let us assume a perfectly rigid inner core ($\tau=\infty$), and no coupling at the CMB (${\cal F}_{m}(s)=0$). Let us further assume that the fluid inside the TC rotates as a single rigid body with angular velocity $\Omega_{f}^{tc}$. Taking the time-derivatives of eqs. (1a), (1b) and (1c), and using eq. (3), gives
$C_{m}\frac{d^{2}\Omega_{m}}{dt^{2}}$ $=\bar{\Gamma}(\Omega_{i}-\Omega_{m})\,,$ (7a)
$C_{i}\frac{d^{2}\Omega_{i}}{dt^{2}}$ $=-\bar{\Gamma}(\Omega_{i}-\Omega_{m})+r_{i}\,\overline{\cal F}_{i}\frac{d}{dt}(\Omega_{f}^{tc}-\Omega_{i})\,,$ (7b)
@@
where $C_{f}^{tc}$ is the moment of inertia of the fluid inside the TC and ¯$\overline{\cal F}_{i}$ is the integrated average of ${\cal F}_{i}(s)$ over the ICB. This forms a simple system of axial oscillations between $\Omega_{m}$, $\Omega_{i}$ and $\Omega_{f}^{tc}$, with a restoring gravitational torque between the inner core and mantle and a restoring (and dissipating) EM torque between the inner core and the fluid inside the TC. Note that the fluid in the ROTC is not involved and assumed to remain at rest. The natural modes of oscillations of this system include the MICG mode, whose frequency depends on the strength of $\overline{\Gamma}$ and ¯$\overline{\cal F}_{i}$.
@@
Let us consider two end-member scenarios for ¯$\overline{\cal F}_{i}$. For no EM coupling at the ICB (¯$\overline{\cal F}_{i}=0$), the fluid inside the TC is decoupled and not involved in the MICG mode. The frequency and eigenvector of the MICG mode are (dumberry10b),
@@
$\omega_{micg}=\sqrt{\frac{\bar{\Gamma}(C_{m}+C_{i})}{C_{m}C_{i}}}\,,\hskip 28.45274pt\left[\Omega_{m},\Omega_{i}\right]=\left[1,-\frac{C_{m}}{C_{i}}\right]\,.$ (8)
@@
The oscillation (and angular momentum exchange) is between the mantle and inner core, with no involvement of any parts of the fluid core. As an example, taking $\overline{\Gamma}=3\times 10^{20}$ N m, and $C_{m}$ and $C_{i}$ given Table 1, this gives a MICG period of 2.8 years. If instead ¯$\overline{\cal F}_{i}$ is very large, the differential rotation between $\Omega_{f}^{tc}$ and $\Omega_{i}$ must remain very small (over the period of the MICG mode), and $\Omega_{f}^{tc}$ and $\Omega_{i}$ are locked into a common rotation. Substituting the EM coupling term of eq. (7c) into eq. (7b), and assuming $\Omega_{f}^{tc}\approx\Omega_{i}$, the frequency and eigenvector of the MICG mode are now

### v2

@@
Abstract The mantle-inner core gravitational (MICG) mode is the free mode of axial oscillation between the mantle and inner core, sustained by the gravitational torque between their degree 2 order 2 density structures. As part of this mode, the tangent cylinder (TC) is entrained to move jointly with the inner core, and the oscillations of the TC launch Alfvén waves propagating in the region outside the TC. Here, we investigate how the MICG is altered by the strength of the internal magnetic field in the core which controls the travelling speed of Alfvén waves. We show that the MICG mode remains a distinct normal mode of oscillation of the core-mantle system only when Alfvén waves are attenuated before they traverse the width of the fluid core. For an internal magnetic field strength of a few mT, as we expect in Earth’s core, Alfvén waves can readily traverse the width of the core, and the MICG mode is absorbed into the spectrum of torsional oscillation (TO) modes. The MICG period retains a dynamical influence, acting as a point of resonance for TO modes, and marking the transition from a TO mode that is weakly impacted by gravitational coupling to one in which the motion of the TC (including the inner core) is strongly restricted. Our results imply that the observed 6-year periodic signal in the length of day cannot be interpreted as the signature of the MICG mode and must instead be caused by TO modes, or more generally, by the propagation of Alfvén waves.
{summary}
Non-technical summary The Earth’s mantle is not uniform but contains continental size density anomalies. The topography of the solid inner core (located at the centre of our planet) is aligned on average with these mantle density anomalies. The axial gravitational restoring torque between these mantle and inner core structures – when they are rotated out of longitudinal alignment – sustains a natural mode of oscillation termed the mantle-inner core gravitational (MICG) mode. Previous studies have suggested that the MICG mode may explain a periodic signal of 6 years observed in the changes in the length of day (LOD). Here, we show that when the magnetic field inside the fluid core is high (of the order of a few mT), the nature of the MICG changes; it is no longer a distinct, independent mode of oscillation of the core-mantle system. Instead, the MICG mode merges into the set of modes of torsional oscillations (TO), which consist in the longitudinal oscillations of nested cylindrical surfaces within the fluid core maintained by the magnetic tension between them (oscillations that are also known as Alfvén waves). The inner core can still oscillate with respect to the mantle, but it is due to its entrainment as part of the TO modes. These results have important implications for the interpretation of the 6 year LOD signal, which cannot be caused by the MICG mode, but must instead be caused by the TO modes, or more generally, by the propagation of Alfvén waves in the fluid core.
@@ 1 Introduction
The mantle-inner core gravitational (MICG) mode of axial oscillation owes its existence to aspherical mass anomalies within the mantle, including those caused by the topography at surfaces of density contrast [7]. The largest of these occur at spherical harmonic degree 2 order 2 [16, 46, e.g.], warping equipotential surfaces everywhere inside the Earth (including the surface geoid) into ellipsoid shapes elongated in the equatorial plane. Within the core, surfaces of constant density coincide with these elliptically shaped equipotential surfaces in order to maintain hydrostatic equilibrium. Provided the inner core deforms viscously on a timescale shorter than the typical changes in the gravitational potential imposed by the mantle, the inner core boundary (ICB) should also match an equipotential surface. Its shape should then be elliptical in the plane of the equator (i.e. equatorially elliptical) and longitudinally aligned with the leading order mantle mass anomalies [7]. An axial rotation of the inner core out of this longitudinal alignment, driven for example by magnetic stresses at the ICB [9, 2, e.g.], leads to a gravitational torque acting to restore the equilibrium state of minimum gravitational potential energy [7]. This gravitational torque can thus sustain a natural mode of axial oscillation between the inner core and the mantle, referred to as the MICG mode [39].
@@
The period of the MICG mode depends on the gravitational strength factor, $\overline{\Gamma}$, capturing the amplitude of the gravitational torque for a given small longitudinal misalignment between the inner core and mantle [6]. In turn, $\overline{\Gamma}$ depends on the amplitude of the degree 2 order 2 geoid at the core-mantle boundary (CMB), which itself depends on estimates of the mantle density anomalies, including that caused by the topography of the CMB [7, 15, 12, 45]. The period of the MICG mode also depends on whether the region of the fluid core inside the axial cylinder that encloses the inner core – the so-called tangent cylinder (TC) – is entrained into co-rotation with the inner core, which is expected given the strong electromagnetic coupling at the ICB [6, 37].
@@
Estimates of the period of the MICG mode can be derived by computing predictions of $\overline{\Gamma}$ on the basis of viscous mantle flow models. The latter are built from density anomalies within the mantle inferred from seismic tomography and the subsequent viscous deformation of surfaces of density discontinuities that they cause [28, 22, 16, e.g.]. Based on the mantle flow models available at the time, [7] estimated a MICG period of approximately 2 to 3 years. The more recent survey of [15] suggests a longer period in the range 7 to 18 years. The range of uncertainty reflects our limited knowledge of both the amplitude of the internal mantle density anomalies [23, 36, primarily those of degree 2, order 2 associated with Large Low Seismic Velocity provinces in the lower mantle, e.g.] and the viscosity of the lowermost mantle.
@@
An alternate approach is to seek the signature of the MICG mode in the observed changes in the mantle rotation rate measured as changes in the length of day (LOD). Indeed, an observed quasi-periodic LOD signal with an amplitude of $\sim 0.13$ ms and a period of $\sim 5.9$ years [1, 30, 13] has been linked with the MICG mode [39, 12, 17, 18, 45]. The magnitude of $\overline{\Gamma}$ required to match this period provides then a direct measure of the equatorial ellipticities of the geoid at the CMB and of the ICB topography [15, 12, 45].
@@
This simplified picture of the MICG mode neglects the dynamical influence of the fluid core in the region outside the tangent cylinder (ROTC) or assumes that it plays a passive role. At periods relevant to the MICG mode, axisymmetric azimuthal (zonal) flows are expected to be invariant in the direction of the rotation axis and take the form of rigidly rotating nested cylindrical surfaces [32]. A differential motion between adjacent cylinders shears the cylindrically radial ($B_{s}$) magnetic field lines that threads their surfaces. This creates an azimuthal restoring force between cylinders and allows for the propagation of Alfvén waves in the cylindrically radial direction [4].
The normal modes of axial oscillations between the cylinders formed by the superposition of these Alfvén waves are referred to as torsional oscillations [4, TO,] and their periods depend the strength of $B_{s}$ within the core. These TO modes include the motion of the TC, so they are necessarily coupled to the MICG mode. The prevalent view at the time the earliest studies on the MICG mode were published [7, 6, 37, 39, e.g.] was that the gravest mode of TO had a period of 60-80 years, implying a $B_{s}$-field strength of 0.2-0.3 mT within the core [51]. Under this scenario, the MICG period is distinctly shorter than that of the fundamental TO mode. Consequently, the MICG mode provides a resonant band of amplification to higher TO harmonics that have periods in the vicinity of the MICG period [37]. The oscillating motion of the inner core, entrained by the TO modes, exerts a gravitational torque on the mantle leading to distinct peaks in the LOD at periods close to the MICG mode. Even though the ROTC is dynamically involved, it carries little angular momentum and the primary angular momentum balance for these amplified TO modes is between the mantle and the whole of the TC, just as is the case for a pure MICG mode. The 5.9 yr oscillation in the LOD may still be tied to MICG period and provides a constraint on $\overline{\Gamma}$ [39].
@@
The general consensus on the magnetic field strength within the core has since shifted. Numerical models of the dynamo suggest that the field deep inside the core is likely stronger by a factor 10 than at the CMB, suggesting an internal field strength of 3-4 mT [14, 3, e.g.]. With such a large field strength, Alfvén waves propagate through the entire width of the core in only a few years and the periods of the gravest TO modes are expected to be in the subdecadal range. Indeed, [25] showed that the angular momentum carried by propagating Alfvén waves matches that required to explain the observed periodic $5.9$-yr LOD signal. This indicates that the mantle exchanges its angular momentum with the whole of the core rather than with the TC alone, in contrast with an idealized MICG mode. Furthermore, in this scenario, the 5.9 yr period is instead tied to the period of the fundamental mode of TO, not the strength of the gravitational torque, even though the gravitational torque may still provide the means by which angular momentum is transferred to the mantle.
@@
In a strong $B_{s}$-field regime, Alfvén waves travel through the width of the core in a shorter time than the MICG period. Here, our goal is to investigate how the MICG mode is altered by such fast propagating Alfvén waves. We shall show that when Alfvén waves can propagate through the ROTC before they get attenuated, the MICG mode is no longer an independent mode of the coupled core-mantle system. Instead, it gets absorbed into the spectrum of TO modes, although the latter are affected by gravitational coupling.
@@ 2.1 Angular momentum balance
The model of internal coupling between the mantle, fluid core and inner core that we use follows broadly that presented in several studies, including in [6, 8, 37, 38, 20, 21]. We assume a spherical model of Earth’s interior comprised of an inner core (radius $r_{i}$), a fluid core (outer radius $r_{f}$) and mantle. We adopt a simplified model of flow in the fluid core restricted to its azimuthal component $v_{\phi}$ and assumed to be axisymmetric and invariant along the direction of the rotation axis. Accordingly, we use cylindrical coordinates ($s$, $\phi$, $z$), and we divide the fluid core in a set of $N_{s}$ discrete cylindrical surfaces with radius $s$ that intersect the CMB at axial position $z_{f}=\sqrt{r_{f}^{2}-s^{2}}$ in the Northern hemisphere. Cylinders inside the TC ($s<r_{i}$) intersect the ICB at $z_{i}=\sqrt{r_{i}^{2}-s^{2}}$; we set $z_{i}=0$ for cylinders outside the TC. The azimuthal velocity of each cylinder is set equal to $v_{\phi}=s\Omega_{f}$, where $\Omega_{f}\equiv\Omega_{f}(s)$ is the angular velocity.
@@
is the axial moment of inertia density of cylinders within the fluid core, where $\rho$ is the density (assumed uniform). $\Gamma_{cmb}$ and $\Gamma_{icb}$ are the torques from all surface forces acting on the mantle at the CMB and on the inner core at the ICB, respectively. $\overline{\Gamma}$ is the strength of the gravitational torque between the mantle and inner core and $\alpha$ is the (assumed small) longitudinal angle of misalignment between the degree 2 order 2 mantle density field and the long equatorial axis of the ICB topography. $\alpha$ depends on the differential rotation between the inner core and the mantle and on the viscous relaxation time $\tau$} for the ICB topography to return to its equilibrium alignment with the imposed gravitational potential from the mantle; its evolution is given by
@@
We assume that $\Gamma_{icb}$ and $\Gamma_{cmb}$ are both dominated by electromagnetic (EM) forces, and expressed as [8, e.g.]
$\Gamma_{cmb}$ $=\int_{0}^{r_{f}}{\cal F}_{m}(s)\left[\Omega_{f}-\Omega_{m}\right]ds\,$, (4a)
$\Gamma_{icb}$ $=\int_{0}^{r_{i}}{\cal F}_{i}(s)\left[\Omega_{f}-\Omega_{i}\right]ds\,$, (4b)
@@
where $B_{m}^{d}$ and $B_{i}^{d}$ are the polar amplitudes of the axial dipole at the CMB and ICB, respectively, $\langle B_{r,m}\rangle$ is the root mean squared (rms) strength of the total radial field (including the axial dipole) at the CMB, $G_{m}$ is the conductance of the lowermost mantle and $G_{i}$ is a conductance factor at the ICB given by
@@
where $\sigma$ is the conductivity of the fluid core (assumed equal to that of the inner core), $\delta=\sqrt{2/|\omega|\mu\sigma}$ is the magnetic skin depth at oscillation frequency $\omega$, with $\mu$ the magnetic permeability of free space, and $\mbox{sgn}(\omega)=\omega/|\omega|$. The form of ${\cal F}_{m}(s)$ was shown in [20] to be a good match with the EM coupling resulting from a realistic field model at the CMB comprised of a large number of spherical harmonic terms. The field at the ICB is undoubtedly more complex than a simple axial dipole, although the latter is expected to be its dominant contribution, and it is sufficient to capture the first order effect of EM coupling at the ICB.
@@ 2.2 The MICG mode
To illustrate how the MICG mode emerges from the above system, let us assume a perfectly rigid inner core ($\tau=\infty$), and no coupling at the CMB (${\cal F}_{m}(s)=0$). Let us further assume that the fluid in the ROTC is not involved and remains at rest, while the fluid inside the TC rotates as a single rigid body with angular velocity $\Omega_{f}^{tc}$. Taking the time-derivatives of eqs. (1a), (1b) and (1c), and using eq. (3), gives
$C_{m}\frac{d^{2}\Omega_{m}}{dt^{2}}$ $=\overline{\Gamma}(\Omega_{i}-\Omega_{m})\,$, (7a)
$C_{i}\frac{d^{2}\Omega_{i}}{dt^{2}}$ $=-\overline{\Gamma}(\Omega_{i}-\Omega_{m})+r_{i}\,\overline{\cal F}_{i}\frac{d}{dt}(\Omega_{f}^{tc}-\Omega_{i})\,$, (7b)
@@
where $C_{f}^{tc}$ is the moment of inertia of the fluid inside the TC and $\overline{\cal F}_{i}$ is the integrated average of ${\cal F}_{i}(s)$ over the ICB. This forms a simple system of axial oscillations between $\Omega_{m}$, $\Omega_{i}$ and $\Omega_{f}^{tc}$, with a restoring gravitational torque between the inner core and mantle and a restoring (and dissipating) EM torque between the inner core and the fluid inside the TC. The natural modes of oscillations of this system include the MICG mode, whose frequency depends on the strength of $\overline{\Gamma}$ and $\overline{\cal F}_{i}$.
@@
Let us consider two end-member scenarios for $\overline{\cal F}_{i}$. For no EM coupling at the ICB ($\overline{\cal F}_{i}=0$), the fluid inside the TC is decoupled and not involved in the MICG mode. The frequency and eigenvector of the MICG mode are [21, e.g.],
@@
$\omega_{micg}=\sqrt{\frac{\bar{\Gamma}(C_{m}+C_{i})}{C_{m}C_{i}}}\,,\hskip 17.07182pt\left[\Omega_{m},\Omega_{i}\right]=\left[1,-\frac{C_{m}}{C_{i}}\right]\,.$ (8)
@@
The oscillation (and angular momentum exchange) is between the mantle and inner core, with no involvement of the fluid core. As an example, taking $\overline{\Gamma}=3\times 10^{20}$ N m, and $C_{m}$ and $C_{i}$ given in Table 1, this gives a MICG period of 2.8 years. If instead $\overline{\cal F}_{i}$ is very large, the differential rotation between $\Omega_{f}^{tc}$ and $\Omega_{i}$ must remain very small (over the period of the MICG mode), and $\Omega_{f}^{tc}$ and $\Omega_{i}$ are locked into a common rotation. Substituting the EM coupling term of eq. (7c) into eq. (7b), and assuming $\Omega_{f}^{tc}\approx\Omega_{i}$, the frequency and eigenvector of the MICG mode are now
$\omega_{micg}=\sqrt{\frac{\bar{\Gamma}(C_{m}+C_{tc})}{C_{m}C_{tc}}}\,,\hskip 17.07182pt\left[\Omega_{m},\Omega_{i}\right]=\left[1,-\frac{C_{m}}{C_{tc}}\right]\,$, (9)
where $C_{tc}=C_{i}+C_{f}^{tc}$. Even though the gravitational torque remains between the mantle and inner core, the oscillation in this case is between the mantle and the whole of the TC (including the inner core). The larger moment of inertia of the whole of the TC, compared to the inner core alone, reduces the frequency of the MICG and lengthens its period. Using $\overline{\Gamma}=3\times 10^{20}$ N m as above, the period is now 5.9 years, approximately twice as long. If $\overline{\cal F}_{i}$ is somewhere between these two extremes, with an amplitude such that the time it takes for $\Omega_{i}$ and $\Omega_{f}^{tc}$ to be brought back into co-rotation is of the order a few years, then the MICG mode is more complex, with each of $\Omega_{m}$, $\Omega_{i}$ and $\Omega_{f}^{tc}$ oscillating at different phases.
At periods of a few years, and assuming a radial magnetic field strength at the ICB of the order of 1 mT or larger, EM coupling is sufficiently strong to prevent any large differential rotation at the ICB [27]. A rotation of the inner core will tend to entrain the fluid in contact with the ICB into co-rotation. This zonal motion is communicated by the (strong) Coriolis force to the whole of the fluid column above and below the inner core, forcing it to rotate rigidly with it in accordance with the Proudman-Taylor constraint [32]. Hence, on dynamical grounds, and for the expected strength of the magnetic field at the ICB, we expect that the MICG mode involves the whole of the TC, not just the inner core.
This simple presentation of the MICG mode is idealized in several ways. First, although $\overline{\cal F}_{i}$ is most likely strong enough to lock the fluid cylinders inside the TC to the rotational motion of the inner core, some degree of differential rotation can occur near the outer edge of the TC [6, 8, e.g.]. Second, we have neglected the global elastic deformations that occur in response to the rotational displacement of the inner core out of its equilibrium longitudinal alignment. Taking these into account would lengthen the MICG period, though not by a large fraction. Third, if the inner core relaxes viscously over a relatively short time (approaching the MICG period), this would attenuate the mode significantly.
Perhaps the most important dynamical omission from this idealized picture though is the role played by the fluid in the ROTC. The latter has been considered in a number of studies [37, 38, 10, 39, e.g.], although under the assumption that the magnetic field permeating the entire core is relatively weak, with a strength of the order of 0.3 to 0.4 mT. The primary motivation for our study is precisely to reassess how the dynamics of the ROTC affects the MICG mode when the magnetic field inside the core is much stronger.
2.3 Torsional oscillation modes
The dynamical equation that governs the evolution of $\Omega_{f}$ is obtained by integrating the azimuthal component of the momentum equation over a cylindrical surface in the fluid core. For this specific geometry, the Coriolis acceleration term vanishes [47, 4, e.g.], giving